# CAN-STRESS: A Real-World Multimodal Dataset for Understanding Cannabis Use, Stress, and Physiological Responses

### v1

@@ Abstract
Coping with stress is one of the most frequently cited reasons for chronic cannabis use. Therefore, it is hypothesized that cannabis users exhibit distinct physiological stress responses compared to non-users, and these differences would be more pronounced during moments of consumption. However, there is a scarcity of publicly available datasets that allow such hypotheses to be tested in real-world environments. This paper introduces a dataset named CAN-STRESS, collected using Empatica E4 wristbands. The dataset includes physiological measurements such as skin conductance, heart rate, and skin temperature from 82 participants (39 cannabis users and 43 non-users) as they went about their daily lives. Additionally, the dataset includes self-reported surveys where participants documented moments of cannabis consumption, exercise, and rated their perceived stress levels during those moments. In this paper, we publicly release the CAN-STRESS dataset, which we believe serves as a highly reliable resource for examining the impact of cannabis on stress and its associated physiological markers.
@@ I Introduction
Due to its widespread consumption and the increasing legalization and decriminalization in many regions, studying cannabis use and its implications is essential [1]. Understanding the physiological and psychological effects of cannabis is crucial for informing policy, healthcare, and individual decision-making. Furthermore, differentiating between the physiological responses of cannabis users and non-users offers insights into how chronic use may alter stress regulation, a frequently cited reason for cannabis consumption [2]. However, most existing studies are conducted in controlled laboratory environments, where ecological validity is constrained [3]. These settings fail to capture the nuanced, real-world conditions under which individuals consume cannabis, including the interplay of stressors and other contextual factors. Bridging this gap through field-based research is imperative to develop a comprehensive understanding of the effects of cannabis’ under realistic settings and enhance the applicability of findings to broader populations and real-world scenarios.
@@
Recently, wearable devices have shown great potential for studying individuals as they go about their daily routines[4, 5]. Many can collect various data modalities simultaneously as they are equipped with multiple sensors [6, 7]. Their small size and user-friendly design minimize disruption to participants’ everyday activities, thereby reducing the introduction of confounding variables during subsequent analyses[8, 9, 10]. Moreover, their wireless connectivity provides researchers with real-time control, which closely approximates the level of oversight achievable in laboratory settings[11, 12]. Additionally, their built-in processing units can handle advanced signal processing algorithms [11, 13, 14, 15].
To advance research on cannabis consumption and its physiological effects, this paper presents CAN-STRESS, a multimodal dataset gathered using wearable wristbands from 82 participants during a full day as they engaged in their usual daily activities. The data collection process, including participant recruitment, study protocols, and the use of self-reported questionnaires to document daily activities such as cannabis use, sleep, exercise, and perceived stress levels, has been thoroughly detailed in [16].
This dataset is one of the largest publicly available resources capturing the physiological characteristics of cannabis users in real-world settings. By including both cannabis users and non-users, it facilitates balanced comparisons and robust analysis. CAN-STRESS offers a range of modalities, including Electrodermal Activity (EDA), Heart Rate (HR), body temperature, and accelerometer readings, with synchronized and high-quality data that address the limitations of previous laboratory-based studies. The dataset is highly suitable for diverse research applications, such as exploring stress biomarkers, analyzing the physiological effects of cannabis consumption [16], and developing advanced signal processing algorithms[17].
In the next section, we briefly discuss the process of collecting the data. We then outline the various components of the dataset. In section four, we highlight key characteristics observed within the dataset. We then present some applications and preliminary results derived from our dataset in section five. Finally, the conclusion is provided in section six.
II Data Collection
Participants for this study were recruited from the Palouse region of Washington State over a period spanning February 2022 to November 2023. To qualify, individuals needed to be at least 21 years old, proficient in English, and in possession of a smartphone. Individuals were excluded if they had a history of neurological conditions, such as head injuries or seizures, or severe psychiatric disorders like psychosis, autism, or bipolar I disorder. Furthermore, participants who reported heavy alcohol use—defined as consuming more than four drinks on more than four occasions per week—were excluded, as were those who had used illicit drugs, nicotine, or tobacco products in the past three months or were currently taking corticosteroid medications. Cannabis users in the study were required to report daily or near-daily consumption (at least four times per week) over the past year, while non-users had to have consumed cannabis fewer than ten times in their lifetime, with no use in the previous year.
Recruitment was carried out using a combination of online advertisements and physical postings in local businesses, including cannabis dispensaries, coffee shops, and laundromats. These advertisements featured a URL and a scannable QR code, both of which directed interested individuals to a brief online screening survey. At the end of the survey, eligible participants were invited to provide their email addresses if they wished to be contacted for the study. Once eligibility was confirmed by a researcher, participants were contacted to schedule an initial Zoom meeting. During this meeting, the researcher explained the study’s procedures in detail and obtained informed consent. Following the meeting, study materials were delivered to the participants’ homes, including the Empatica E4 [12], a medical-grade wearable device (figure 2), along with detailed written instructions. Participants were asked to wear the wristband on their non-dominant wrist continuously for a 24-hour period, beginning as soon as they woke up on the designated day of data collection and ending the following morning, and they were instructed not to use alcohol on the study day to avoid confounds. For additional details regarding the collection of CAN-STRESS, please refer to [16].
Figure 2: Empatica E4 wristband
III Data Modalities
Figure 3: Comparison of Dataset Features Between Users and Non-Users
@@ III Data Modalities
Our dataset consists of two major modalities: a self-reported questionnaire and multimodal physiological data collected through the E4 wearable wristband. Together, these modalities provide a comprehensive view of participants’ activities, physiological responses, and subjective experiences. Aligning and integrating the two sources of the dataset allows researchers to identify and analyze the correlations between different events, e.g. cannabis consumption vs exercise.
III-A Self-Reported Questionnaire
@@
  • Cannabis Use: Participants documented the start and end times of their cannabis consumption.
@@
  • Stress Ratings: At these moments (e.g., cannabis use, exercise, sleep), participants rated their perceived stress on a scale of 1 to 10, which provided subjective insights into their mental state throughout the day.
III-B Physiological Data from the Wristband
@@
  • Blood Volume Pulse (BVP): Measured at 64 Hz, BVP provides raw data for calculating heart rate and interbeat intervals that offers insights into cardiovascular dynamics [18].
@@
  • Interbeat Interval (IBI): Derived from BVP, IBI represents the time between successive heartbeats and is critical for heart rate variability analysis [20].
@@
These physiological signals provide high-resolution, multivariate data that capture participants’ physical states and responses in real time. By combining the self-reported data with wristband signals, researchers can investigate the interplay between subjective experiences (e.g., stress ratings) and objective physiological responses (e.g., changes in HR or EDA) across various activities and contexts. [21]
IV Data Features
@@
To provide a comprehensive overview of the dataset, we present key statistical features that highlight differences between cannabis users and non-users, as well as physiological variations associated with self-reported stress levels. These analyses help characterize both baseline physiological patterns and stress-related responses.
IV-A Group-Level Comparisons: Users vs. Non-Users
This subsection presents key statistical features comparing cannabis users and non-users. These features include the average recording duration, self-reported stress ratings, sleep times, and physiological signals such as EDA and HR. Additionally, we include EDA peaks per minute (PPM), a well-established metric closely associated with stress, which provides valuable insights into stress-related physiological responses. These statistics offer insights into the typical physiological and behavioral differences between the two groups and serve as a foundation for further analyses.
Table I provides a detailed comparison of cannabis users and non-users across various dimensions. For each group, we report the number of participants, the average recording time (hours), the average self-reported stress rating (1–10 scale), the average sleep duration (hours/day), the mean EDA value (microsiemens), the mean EDA PPM, and the mean HR (beats per minute). Standard deviations are also included to illustrate the variability within each group. These features are further visualized in Figure 3, which presents boxplots for each metric. This figure highlights key trends such as higher stress levels, EDA values, and PPM in users, alongside longer sleep durations in non-users.
TABLE I: Summary of Dataset Features for Users and Non-Users
Feature Users Non-Users
Number of Subjects 39 43
Average Recording Time (hrs) $22.09\pm 4.54$ $21.40\pm 5.45$
Average Stress Rating (1–10) $5.27\pm 2.48$ $4.90\pm 2.47$
Average Time Slept (hrs/day) $7.97\pm 0.83$ $8.17\pm 1.08$
Average EDA Value ($\mu$S) $1.54\pm 1.34$ $1.18\pm 1.03$
Average EDA Peaks (PPM) $1.03\pm 0.34$ $0.97\pm 0.31$
Average Heart Rate (BPM) $91.68\pm 7.52$ $88.77\pm 5.42$
IV-B Physiological Features Around Stress Events
In addition to examining differences between cannabis users and non-users, we analyzed how physiological signals change in response to self-reported stress events. Specifically, we investigated fluctuations in EDA and HR surrounding these moments. To systematically assess these changes, we stratified self-reported stress ratings into four categories: no stress (0–1), low stress (2–4), medium stress (5–7), and high stress (8–9).
For each stress-labeled moment in the dataset, we extracted 60-second segments of EDA and HR data recorded during that period and calculated the average and the standard deviation of the signals to characterize their behavior. As expected, both EDA and HR values tend to be lower in moments associated with lower reported stress levels, while higher stress levels correspond to an increase in physiological activation. Figures 4 and 5 and table II visualize these differences, showing the distribution of EDA and HR across the four stress categories. As you can see, both these signals fluctuate in a much higher range for hight rates of stress. These findings support the expected relationship between physiological arousal and stress intensity and provide a basis for further modeling and predictive analysis.
TABLE II: Mean and Standard Deviation of Physiological Signals Across Stress Levels
Signal No Stress Low Stress Mild Stress High Stress
EDA $0.85\pm 0.11$ $0.73\pm 0.06$ $1.12\pm 0.07$ $1.79\pm 0.08$
HR $88.36\pm 3.89$ $85.23\pm 2.79$ $87.58\pm 2.31$ $89.55\pm 4.63$
Figure 4: Averaged EDA signals recorded over 60-second intervals surrounding moments rated as no, low, mild, and high stress. As shown, EDA tends to be higher during moments of greater reported stress. The shaded regions around the signal represent the standard deviation which illustrates variability across subjects.
Figure 5: Averaged HR signals recorded over 60-second intervals surrounding moments rated as no, low, mild, and high stress. HR increases during high-stress moments, with a more rapid rise compared to lower stress levels. The shaded regions around the signal indicate the std of the signal, reflecting participant variability in cardiovascular responses.
V Application and Use Cases
The CAN-STRESS dataset has proven to be an invaluable resource for advancing research in physiological monitoring, cannabis consumption detection, and stress-related behaviors in real-world settings. Below, we discuss three studies that utilized this dataset and highlight how its multimodal design and real-world applicability contributed to their findings.
The first study [10] leveraged the CAN-STRESS dataset to develop a system for identifying cannabis consumption sessions in naturalistic environments. The authors of [10] focused on using Electrodermal Activity (EDA), a key physiological marker of stress, to model cannabis use detection as a supervised learning problem. By training a neural network on data collected from wearable sensors, the authors demonstrated the feasibility of detecting cannabis use with high accuracy in uncontrolled settings. Additionally, the dataset’s participant diversity and real-world conditions allowed the team to apply transfer learning techniques, which significantly improved the system’s performance for individual users. This dataset enabled the development of a robust and scalable detection framework that would have been impossible to design using laboratory-collected data alone, given its limited variability and constrained scenarios.
In another study [22], the CAN-STRESS dataset was utilized to explore the potential of self-supervised learning for detecting cannabis consumption with limited labeled data. The researchers introduced the CUDLE (Cannabis Use Detection with Label Efficiency) framework, which first used data augmentation techniques and a contrastive learning approach to train models on unlabeled wearable sensor data. The dataset’s extensive physiological measurements and user-reported moments of consumption provided the ideal foundation for validating this novel approach. By combining real-world data with self-supervised pretext tasks, the authors showcased the effectiveness of the dataset in enabling high-performance detection models while significantly reducing reliance on large-scale labeled datasets. This research highlights how CAN-STRESS supports innovation in computational health by addressing challenges unique to free-living environments.
Finally, the CAN-STRESS dataset was instrumental in a study investigating disruptions in diurnal stress rhythms among cannabis users[16]. The wearable device’s ability to capture continuous physiological signals, such as EDA and heart rate variability, alongside self-reported stress ratings, allowed researchers to assess stress-related biomarkers in real-world contexts. These naturalistic conditions were critical in revealing patterns like blunted cortisol awakening responses and elevated afternoon cortisol levels in cannabis users, which traditional lab-based studies often fail to capture. A key finding of [16] is that cortisol levels decreased after cannabis use in people’s natural environments as lab studies have found that cortisol increases after cannabis use in the lab. This finding is more consistent with the fact that many cannabis users use cannabis to decrease stress. The dataset’s inclusion of both users and non-users provided a balanced framework for comparison and enabled a more comprehensive understanding of how chronic cannabis use may affect stress regulation in daily life.
These studies demonstrate how the CAN-STRESS dataset’s real-world applicability, ecologically valid design, and rich labeling have facilitated groundbreaking research across a range of disciplines. Its contributions underscore the value of field-based datasets in advancing our understanding of cannabis consumption, stress, and physiological monitoring.
VI Conclusion

### v2

# Ryan McLaughlin2, Michael J. Cleveland2, Hassan Ghasemzadeh1

@@ Abstract
Coping with stress is one of the most frequently cited reasons for chronic cannabis use. Therefore, it is hypothesized that cannabis users exhibit distinct physiological stress responses compared to non-users, and that these differences may be especially pronounced during moments of cannabis consumption. However, there is a scarcity of publicly available datasets that allow such hypotheses to be tested under real-world conditions. This paper introduces a dataset named CAN-STRESS, collected using Empatica E4 wristbands. The dataset includes multimodal physiological measurements (such as skin conductance, heart rate, and skin temperature) from 82 participants (39 cannabis users and 43 non-users) as they went about their daily routines. In addition to sensor data, participants provided self-reported survey responses that included perceived stress ratings and timestamps of key daily events such as cannabis use, physical activity, and sleep. To demonstrate the utility of the dataset for downstream applications, we present a preliminary machine learning task aimed at classifying cannabis users versus non-users based on physiological features. Our model achieves a classification accuracy of approximately 96% and an f1-score of around 98%. An analysis of feature importance using SHAP values revealed that electrodermal activity and heart rate metrics were the most influential predictors, consistent with their established roles in stress detection. We publicly release the CAN-STRESS dataset, which we believe serves as a reliable and rich resource for studying the physiological correlates of cannabis use and stress in naturalistic settings.
@@ I Introduction
Due to its widespread consumption and the increasing legalization and decriminalization in many regions, studying cannabis use and its implications is essential [12]. Understanding the physiological and psychological effects of cannabis is crucial for informing policy, healthcare, and individual decision-making. Furthermore, differentiating between the physiological responses of cannabis users and non-users offers insights into how chronic use may alter stress regulation, a frequently cited reason for cannabis consumption [1]. However, most existing studies are conducted in controlled laboratory environments, where ecological validity is constrained [8]. These settings fail to capture the nuanced, real-world conditions under which individuals consume cannabis, including the interplay of stressors and other contextual factors. Bridging this gap through field-based research is imperative to develop a comprehensive understanding of cannabis effects under realistic settings and enhance the applicability of findings to broader populations and real-world scenarios.
@@
The CAN-STRESS dataset was developed to enable the study of cannabis use and stress in ecologically valid, real-world conditions. It consists of multimodal data from 82 participants who wore Empatica E4 wristbands [5] for 24 hours while engaging in daily activities. The dataset includes electrodermal activity (EDA), heart rate (HR), body temperature, and accelerometer data, synchronized with self-reported logs of cannabis use, sleep, exercise, and perceived stress levels. By including both frequent cannabis users and non-users, CAN-STRESS provides a balanced framework for comparative analyses and is one of the largest publicly available datasets capturing the physiological characteristics of cannabis users in natural environments.
Participants were recruited with strict inclusion and exclusion criteria. Cannabis users reported daily or near-daily use ($\geq$4 times per week for at least one year), while non-users reported fewer than 10 lifetime uses and none in the past year. Individuals were excluded if they had neurological disorders, psychosis, autism, bipolar I, heavy alcohol consumption, recent use of illicit drugs, nicotine, or corticosteroid medications. Eligible participants were at least 21 years old, fluent in English, and smartphone owners. The study was reviewed and approved by the Washington State University Institutional Review Board, and informed consent was obtained prior to participation. To protect participant privacy, all wearable and survey data were anonymized and securely stored, with study materials collected directly by researchers to minimize risk of exposure.
Several studies have already leveraged CAN-STRESS to advance cannabis research.  [3] demonstrated that EDA signals can be used to detect cannabis consumption episodes in naturalistic environments.  [2] introduced the CUDLE framework, showing that self-supervised learning can efficiently detect cannabis use with limited labeled data. A third work examined stress regulation, revealing disrupted diurnal stress rhythms among cannabis users and showing that cannabis decreased stress in daily life, in contrast to laboratory findings [4]. Together, these works highlight the dataset’s value in enabling ecologically valid and computationally innovative research.
Building on this foundation, the present paper contributes (1) a detailed description of the dataset, (2) new analyses comparing physiological features of users vs. non-users, and (3) a baseline machine learning pipeline that distinguishes users from non-users using individual-level features. These contributions aim to broaden access to the dataset and establish benchmarks for future computational health research.
II Data Modalities
Figure 2: Comparison of dataset features between cannabis users (n=39) and non-users (n=43). Each boxplot represents participant-level summary values: (1) total recording duration (hours), (2) mean self-reported stress rating across the day, (3) mean electrodermal activity (EDA, µS), and (4) mean heart rate (bpm)
@@ II Data Modalities
The CAN-STRESS dataset includes two primary modalities: a self-reported questionnaire and multimodal physiological data collected through the E4 wearable wristband. Together, these modalities provide a comprehensive view of participants’ activities, physiological responses, and subjective experiences. Aligning and integrating the two sources of the dataset allows researchers to identify and analyze the correlations between different events, e.g., cannabis consumption and exercise.
The dataset is organized in a straightforward structure to facilitate use by other researchers. All files are arranged under a root directory labeled CAN-STRESS/, with a subfolder for each anonymized participant identifier. Within each participant’s folder, the physiological modalities are stored as individual .csv files (e.g., ACC.csv, EDA.csv, BVP.csv), each containing time-stamped recordings sampled at their respective rates. To aid interpretation, each participant folder also includes an info.txt file describing the contents, units, and sampling rates of the CSV files. The self-reported questionnaire data are compiled across all participants in a single logbook.xlsx file stored at the root level, which documents daily activities such as cannabis use, sleep, exercise, and stress ratings.
II-A Self-Reported Questionnaire
@@
  • Cannabis Use: Participants stated whether or not they were a user, and they also documented the start and end times of their cannabis consumption.
@@
  • Stress Ratings: At key moments (e.g., sleep, exercise, cannabis use), participants rated their perceived stress on a scale from 1 (not at all stressed) to 10 (extremely stressed). This straightforward approach has been widely used in ecological momentary assessment (EMA) studies to capture in-the-moment subjective stress levels [11].
II-B Physiological Data from the Wristband
@@
  • Blood Volume Pulse (BVP): Measured at 64 Hz, BVP provides raw data for calculating heart rate and interbeat intervals that offers insights into cardiovascular dynamics.
@@
  • Interbeat Interval (IBI): Derived from BVP, IBI represents the time between successive heartbeats and is critical for heart rate variability analysis.
@@
These physiological signals provide high-resolution, multivariate data that capture participants’ physical states and responses in real time. By combining the self-reported data with wristband signals, researchers can investigate the interplay between subjective experiences (e.g., stress ratings) and objective physiological responses (e.g., changes in HR or EDA) across various activities and contexts.
III Descriptive Analysis of Cannabis Users and Non-Users
@@
To provide a comprehensive overview of the dataset, we present key statistical features that highlight differences between cannabis users and non-users. These analyses help characterize both baseline physiological patterns and stress-related responses, and serve as a foundation for downstream computational tasks. In particular, we use these features to develop a machine learning model that classifies cannabis users and non-users based on their physiological data.
III-A Group-Level Trends
This subsection presents important statistical distinctions between cannabis users and non-users, covering aspects such as average recording time, self-reported stress levels, sleep duration, and physiological measurements including EDA and heart rate. These variables reflect common behavioral and physiological characteristics within each group and form the basis for additional analyses, including those involving predictive modeling.
Figure 2 presents four boxplots comparing cannabis users (n=39) and non-users (n=43) across key physiological and behavioral dimensions. Recording duration (in hours) reflects the total amount of usable data collected per participant. For all other measures, participant-level means were computed over the 24-hour recording period, including average self-reported stress ratings (on a 1–10 scale), mean electrodermal activity (EDA, in microsiemens), and mean heart rate (in beats per minute). These participant-level summary values were then aggregated within each group and visualized as boxplots to illustrate variability and group-level trends. The results highlight consistent patterns, with cannabis users reporting higher stress ratings and exhibiting elevated EDA and heart rate values compared to non-users.
III-B Machine Learning for User Classification
Building on the group-level feature analysis, we explore the use of machine learning to classify participants as cannabis users or non-users based on their physiological data. The goal of this task is not to develop a production-ready classifier but to demonstrate the feasibility of using wearable data for downstream predictive modeling. The envisioned pipeline takes as input the raw physiological signals collected by the wristband and predicts a participant’s user status. For personalization, we fine-tuned the pre-trained model using 50% of the available windows from each test subject, while reserving the remaining 50% for evaluation. This setup highlights the potential of the dataset for real-world applications that rely on minimal supervision and passive sensing.
The classification model employed a multi-layer perceptron architecture consisting of three hidden layers with 128, 64, and 32 neurons, respectively, followed by a single-neuron output layer for binary classification. Each hidden layer incorporated batch normalization and ReLU activation, with dropout regularization (rate=0.2) applied to mitigate overfitting. The output layer utilized sigmoid activation to produce probability scores. Training and fine-tuning were performed using the Adam optimizer, binary cross-entropy loss, and a learning rate of 0.001.
To prepare the input data for our model, we selected four physiological modalities from the wristband data for our predictive machine learning task: EDA, BVP, ACC, and TEMP. To ensure data relevance and minimize confounding from sleep-related physiological changes, we restricted our analysis to data recorded during participants’ waking hours. The continuous signals were segmented using a sliding window approach with fifteen-minute windows and 50% overlap between consecutive segments [10]. On average, this produced approximately 180 windows per participant, resulting in a total of about 14,600 windows across the dataset. Each window was represented by 31 physiological features extracted from the EDA, BVP, ACC, and TEMP signals.
For feature extraction, we applied modality-specific methods designed to capture the distinct temporal and physiological characteristics of each signal. Using the NeuroKit2 [7] package, we decomposed the EDA signal into tonic and phasic components and extracted features such as mean tonic level, number of skin conductance response peaks, peaks per minute, and the statistical properties of peak amplitudes. From the BVP signal, we derived heart rate variability (HRV) features, including RMSSD, pNN50, and SDNN, in addition to basic statistical measures of heart rate (e.g., mean and standard deviation). For the temperature signal, we computed statistical metrics, including mean, standard deviation, and range, along with the linear slope to capture thermal trends across each window. Accelerometer data was processed by computing the vector magnitude of the triaxial acceleration, from which we extracted features such as mean activity level, proportion of time spent in motion, and axis-specific statistics that characterize both intensity and direction of movement. In total, we extracted 31 features across all modalities, which served as the input to our downstream classification task.
Following feature extraction, our dataset consisted of tabular data, where each row corresponded to a 15-minute window and included 31 physiological features. To address individual differences in baseline physiology, we applied subject-wise standardization, ensuring that each participant’s features had a mean of zero and a standard deviation of one. We employed a leave-one-subject-out strategy to evaluate model generalization to new individuals. In each fold, one participant served as the test subject while the model trained on all remaining participants. To personalize the model for each test subject, we implemented a transfer learning approach: we first trained a base model on the training subjects, then fine-tuned the final layer of the pre-trained model using 50% of the test subject’s data, and evaluated performance on the remaining 50%. This process was repeated across all participants, with each individual serving as the test subject once. Final performance metrics were computed as averages across all folds.
In table I, We report our model’s performance on both the training and test data using four evaluation metrics: accuracy, F1 score , precision, and recall. Accuracy represents the overall proportion of correct predictions. Precision measures the fraction of predicted positive cases that are actually positive, while recall reflects the fraction of actual positive cases that were correctly identified by the model. The F1 score is the harmonic mean of precision and recall, offering a single metric that balances both, particularly useful in scenarios with class imbalance.
TABLE I: Model performance on training and test data (averaged across subjects).
Data Acc Prec Rec F1
Train Data 99.93% ($\pm$0.00) 99.96% ($\pm$0.00) 100.00% ($\pm$0.00) 99.93% ($\pm$0.05)
Test Data 95.96% ($\pm$0.06) 97.82% ($\pm$0.05) 100.00% ($\pm$0.00) 95.92% ($\pm$0.06)
To better understand which signals influenced the model’s predictions, we used SHAP (Shapley Additive Explanations) [6] to analyze feature importance. As shown in the SHAP summary plot (Figure 3), features derived from HR and EDA dominated the top rankings. The most impactful feature was the maximum heart rate (hr_max), followed closely by several EDA-related features, including eda_min, eda_mean, eda_phasic_mean, and eda_max. Heart rate variability metrics such as hrv_sdnn and hrv_rmssd also contributed significantly. This result aligns with the expectation that physiological markers of stress (captured through both heart rate and skin conductance) differ between cannabis users and non-users. Overall, the SHAP results highlight the central role of stress-related signals, particularly EDA and HR, in driving the model’s ability to distinguish between the two groups.
Figure 3: SHAP summary plot showing the top 10 most important features influencing the model’s predictions. Features related to heart rate, heart rate variability, and electrodermal activity contribute most strongly
IV Conclusion